OpenStudy (anonymous):

(dividing parabola's) 2x-7(divided by) 2x^2+7x-39 @Mathematics

jhonyy9 (jhonyy9):

2x2+7x-39= x1,2=(-7+/- sqrt(49+312))/4 =(-7+/- sqrt361)/4=(-7+/-19)/4= -26/4 =-13/2 and 12/4=3 so 2x2 +7x -39=(x-(-13/2))(x-3)=(2x+13)(x-3) so than (2x-7)/(2x+13)(x-3) = 2x/(2x+13)(x-3) - 7/(2x+13)(x-3)